4 ms·
Actually, he meant a 75% chance of winning.. Given the chance of winning a round is p=0.5, the chance of losing is q = 1 - p = 0.5 If you've lost the first ro
by amingilani 9y ago
Actually, he meant a 75% chance of winning..
Given the chance of winning a round is p=0.5, the chance of losing is q = 1 - p = 0.5
If you've lost the first round, the odds of you twice in a row are q = 0.5^(2 rounds) = 0.25
Which means the odds of you winning are p = 1 - q = 1 - 0.25 = 0.75.
- 0x8BADF00D 9y agoRight, but the expected value of the gamble is negative.
- dfinch 9y agoExpected value is 0, right? .75 * 10000 - .25 * 30000